# Can Large Language Models Address Open-Target Stance Detection?

### v6

@@ Abstract
Stance detection (SD) identifies a text’s position towards a target, typically labeled as favor, against, or none. We introduce Open-Target Stance Detection (OTSD), the most realistic task where targets are neither seen during training nor provided as input. We evaluate Large Language Models (LLMs) from GPT, Gemini, Llama, and Mistral families, comparing their performance to the only existing work, Target-Stance Extraction (TSE), which benefits from predefined targets. Unlike TSE, OTSD removes the dependency of a predefined list, making target generation and evaluation more challenging. We also provide a metric for evaluating target quality that correlates well with human judgment. Our experiments reveal that LLMs outperform TSE in target generation, both when the real target is explicitly and not explicitly mentioned in the text. Similarly, LLMs overall surpass TSE in stance detection for both explicit and non-explicit cases. However, LLMs struggle in both target generation and stance detection when the target is not explicit. [^1] Dataset and code are available here: anonymous.4open.science/r/opentarget-5521.
@@ 3.2 Approach
To address the research questions outlined in §1, our approach to OTSD leverages the zero-shot learning capabilities of LLMs. We adopt the "Task Definition" prompting strategy, as demonstrated by Cruickshank and Ng (2023), which either outperforms or performs on par with alternative strategies such as "Chain-of-Thought" (CoT) and "few-shot" across most stance detection datasets. To further validate this, we conduct a preliminary experiment with the CoT strategy (detailed in Appendix J), which reveals that CoT underperforms in TG, and performs comparably to "Task Definition" in SD, reinforcing the early findings. Additionally, using the "Task Definition" strategy without relying on tailored prompts for specific targets or in-context examples ensures the generalizability of OTSD, aligning with the zero-shot nature of this task. This approach first names the task and then provides a concise, self-contained definition of what that task entails—its purpose, inputs, and expected outputs—before asking the model to perform it. We employ two different prompting approaches for TG and SD. The prompts and their design justifications can be found in Appendix A. TG+SD (Two-Step Approach) In this approach, we first generate the target and then sequentially detect the stance of the text towards the generated target, which directs LLMs to focus on one task at a time. TG&SD (Single-Step Joint Approach) This approach unifies TG and SD in a single step using a prompt, enabling the model to better understand the relationship between the text, target, and stance.
@@ 4.2 Evaluation Methods
Human Evaluation: To evaluate the relevance of the generated targets by some models and further validate the BTSD metric, we conducted a small-scale human evaluation on some of the models using 500 randomly selected samples from each dataset, maintaining an equal ratio of the stance labels. Three annotators were asked to assess the relevance of the generated targets to the gold ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related) (details in Appendix I). We observed a solid overall annotator agreement of $\alpha$=0.76 (Krippendorff, 2011) and $\kappa$=0.664 (Fleiss, 1971)(Appendix Table 6).
@@ 5.1 Target Generation (TG)
As shown in Table 2, both human evaluation (HE) and BTSD indicate that LLMs generate higher-quality targets in explicit cases compared to TSE-BestGen, the best model for generating targets not matched to a predefined list. Although TSE-mapped achieves higher HE score in explicit case, its BTSD score is lower. This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix K). In contrast, LLMs mostly generate targets that are highly or partially related to gold targets, resulting in fewer irrelevant outputs and higher BTSD, despite slightly lower average HE scores. When the target is not explicitly mentioned, LLMs produce higher-quality targets than both TSE models. However, the quality drops for LLMs (TSE as well) compared to explicit cases (as per Table 2), with HE scores slightly lower than $0.5$ (refers to partially related target) on average. A manual inspection of the lower quality targets suggest that the lower score in non-explicit cases is due to insufficient context of the text. The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (example from GPT-3.5 in Appendix H.1). When comparing tested proprietary and open LLMs using BTSD and HE, there is no clear advantage in TG (i.e., not all proprietary models outperform open ones, or vice versa). Overall, GPT-4o produces targets of equal or higher quality than other tested proprietary models across the approaches, including explicit and non-explicit target mentions. Among the tested open LLMs, Mistral-large and Llama-3-70B outperform each other under different settings. Comparing between approaches, most models perform better in TG&SD across both datasets, for explicit and non-explicit cases.
@@ 5.2 Stance Detection (SD)
As shown in Table 2, according to the SC (stance classification), LLMs outperform both TSE models in detecting stances in both explicit and non-explicit cases. In general, LLMs perform better in SD when the target is explicitly mentioned in the text compared to when it is not. Our manual analysis shows that in non-explicit cases, the input text lacks sufficient context about the target, leading to lower SC scores. The tested proprietary and open LLMs exhibit competitive performance against each other in both explicit and non-explicit cases for both approaches, with neither showing a clear advantage. When comparing approaches, LLMs generally perform better with the holistic TG&SD strategy in both explicit and non-explicit cases, with the exception of Llama-3-8B, and Mistral-small (for non-explicit). The quality of the generated target (BTSD) shows a stronger correlation with the SC score in explicit cases (avg. $\tau=0.4$) across datasets compared to non-explicit ones (avg. $\tau=0.25$), particularly for VAST and EZSTANCE. Upon manual investigation, we observed that LLMs such GPT-4o may generate highly relevant targets that are closely related to the gold targets but nearly ‘antithetical’ or opposite, leading to a stance reversal in comparison to the gold stance. For example, in VAST, when the gold target is "permit to carry gun", GPT-4o generates "gun control", while GPT-3.5 generates "guns". Although GPT-4o’s target is more closely related to the gold target, it conveys the opposite. GPT-3.5’s target, though broader, aligns with the stance of gold target.
@@ 6 Conclusion
  • This work has highlighted specific challenges of the proposed OTSD task and measured both target quality and stance detection performance. However, our findings underscore the need to improve the measure of stance detection accuracy in OTSD, enabling it to account for the alignment/coherence between the generated target and the stance toward it, given the text and the reference (gold) target.
@@
  • Although we did not claim that the results or findings are true for all LLMs, experimenting with other proprietary LLMs was financially constraining for us.
@@ Appendix F Dataset Samples
Following TSE, our targets are Creationism (Somasundaran and Wiebe, 2010), Gay Rights (Somasundaran and Wiebe, 2010), Climate Change is a Concern (Mohammad et al., 2016a), MeToo Movement (Gautam et al., 2020), Merger of Disney and Fox (Conforti et al., 2020), and Lockdown in New York State (Miao et al., 2020). Samples from TSE and VAST in both explicit and non-explicit settings are given in Table 4 and 5, respectively.
@@ Appendix G SemSim Details
Appendix H Explanation with Samples in TG and SD
H.1 Target Generation
For instance, in the post "I mean, this reads to me like anyone who’s an alcoholic doesn’t go to heaven…", GPT-3.5 generated "Sinful behavior and heaven’s eligibility" as the target instead of the gold target "gay rights".
H.2 Stance Detection
For instance, in the explicit sample text, "Q. from audience: Malcom X said the most disrespected person in America is the black woman, has #MeToo changed this? Gender has been constructed in the image of the white woman, the idea of a level playing field, white women are often in better positions, says @HeidiMirza #CHevents", GPT-3.5 generated the target "Gender roles", which, while related, fails to capture the specific stance-related target "MeToo movement". This mistake reflects the model’s need for broader context to establish the correct link.
Appendix I Human Annotation Process
Three graduate students, specializing in NLP, participated voluntarily as annotators in the evaluation of generated targets. They were chosen for their expertise and familiarity with the stance detection task, enabling them to effectively assess the relevance of the generated targets compared to the golden ones. First, we provided them with an annotation guideline (detailed in §I.1) that included descriptions of the datasets, an overview of the task, and a descriptive scoring scale accompanied by examples. Each annotator then independently scored the targets generated by different LLMs for 500 randomly selected samples (300 explicit and 200 non-explicit). To assess inter-annotator agreement, we calculated Krippendorff’s $\alpha$ (Krippendorff, 2011) and Fleiss’s $\kappa$ (Fleiss, 1971), with the results presented in Table 6. The final score for each generated target was determined based on the majority vote among the annotators. In cases of a tie, the median score was used.
I.1 Annotation Guideline
@@
BTSD (GPT-3.5) 38.47 (38.43) 41.10 (41.67) 47.60 (49.27) 37.88 (39.60) 41.34 (44.25) 48.98 (49.70)
BTSD (GPT-4o) 40.63 (41.55) 41.59 (42.69) 48.12 (49.70) 41.31 (41.92) 42.33 (44.25) 48.76 (50.69)
SC (GPT-3.5) 37.91 (42.68) 47.47 (47.21) 40.12 (39.10) 44.86 (47.61) 47.00 (48.48) 40.74 (40.63)
SC (GPT-4o) 47.36 (44.78) 41.17 (40.77) 43.79 (45.93) 49.14 (46.83) 47.84 (49.38) 45.68 (46.22)
@@
BTSD (GPT-3.5) 30.73 (33.10) 41.54 (38.10) 40.76 (38.78) 32.14 (31.32) 40.38 (38.55) 40.54 (41.89)
BTSD (GPT-4o) 32.72 (35.14) 41.71 (38.55) 41.34 (42.78) 33.44 (36.12) 41.92 (39.84) 43.90 (45.70)
SC (GPT-3.5) 31.88 (32.66) 43.41 (45.54) 37.76 (36.69) 35.46 (33.94) 44.46 (45.80) 38.10 (38.97)
SC (GPT-4o) 36.00 (36.39) 38.27 (39.92) 38.34 (38.50) 38.11 (37.5) 42.81 (43.84) 37.54 (39.26)
Table 7: Preliminary results using the CoT prompting strategy with GPT-3.5 and GPT-4o models on three datasets (TSE, VAST, EZSTANCE), evaluated for both Target Generation (TG) using BTSD and Stance Detection (SD) using SC. Results in the parenthesis are from Table 2 following the "Task Definition" strategy. The results are presented for both Explicit (Ex) and Non-explicit (N-Ex) cases across two approaches: TG+SD and TG&SD.
Appendix J Chain-of-Thought (CoT) Results
@@
Figure 1: Distribution of human evaluation scores for target relevance assessment (0 – Not Related, 0.5 – Partially or Indirectly Related, 1 – Completely Related). The final score for each sample is determined by the majority vote among the three annotators. In cases of a tie, the average score of 0.5 is assigned. The details on human annotator guideline are provided in Appendix I.1, and the annotator agreement is in Appendix Table 6.

### v7

@@ Abstract
Stance detection (SD) identifies a text’s position towards a target, typically labeled as favor, against, or none. We introduce Open-Target Stance Detection (OTSD), the most realistic task where targets are neither seen during training nor provided as input. We evaluate Large Language Models (LLMs) from GPT, Gemini, Llama, and Mistral families, comparing their performance to the only existing work, Target-Stance Extraction (TSE), which benefits from predefined targets. Unlike TSE, OTSD removes the dependency of a predefined list, making target generation and evaluation more challenging. We also provide a metric for evaluating target quality that correlates well with human judgment. Our experiments reveal that LLMs outperform TSE in target generation, both when the real target is explicitly and not explicitly mentioned in the text. Similarly, LLMs overall surpass TSE in stance detection for both explicit and non-explicit cases. However, LLMs struggle in both target generation and stance detection when the target is not explicit. [^1] Dataset and code are available here: https://github.com/AbuUbaida/opentarget.
@@ 3.2 Approach
To address the research questions outlined in §1, our approach to OTSD leverages the zero-shot learning capabilities of LLMs. We adopt the "Task Definition" prompting strategy, as demonstrated by Cruickshank and Ng (2023), which either outperforms or performs on par with alternative strategies such as "Chain-of-Thought" (CoT) and "few-shot" across most stance detection datasets. To further validate this, we conduct a preliminary experiment with the CoT strategy (detailed in Appendix I), which reveals that CoT underperforms in TG, and performs comparably to "Task Definition" in SD, reinforcing the early findings. Additionally, using the "Task Definition" strategy without relying on tailored prompts for specific targets or in-context examples ensures the generalizability of OTSD, aligning with the zero-shot nature of this task. This approach first names the task and then provides a concise, self-contained definition of what that task entails—its purpose, inputs, and expected outputs—before asking the model to perform it. We employ two different prompting approaches for TG and SD. The prompts and their design justifications can be found in Appendix A. TG+SD (Two-Step Approach) In this approach, we first generate the target and then sequentially detect the stance of the text towards the generated target, which directs LLMs to focus on one task at a time. TG&SD (Single-Step Joint Approach) This approach unifies TG and SD in a single step using a prompt, enabling the model to better understand the relationship between the text, target, and stance.
@@ 4.2 Evaluation Methods
Human Evaluation: To evaluate the relevance of the generated targets by some models and further validate the BTSD metric, we conducted a small-scale human evaluation on some of the models using 500 randomly selected samples from each dataset, maintaining an equal ratio of the stance labels. Three annotators were asked to assess the relevance of the generated targets to the gold ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related) (details in Appendix H). We observed a solid overall annotator agreement of $\alpha$=0.76 (Krippendorff, 2011) and $\kappa$=0.664 (Fleiss, 1971)(Appendix Table 6).
@@ 5.1 Target Generation (TG)
As shown in Table 2, both human evaluation (HE) and BTSD indicate that LLMs generate higher-quality targets in explicit cases compared to TSE-BestGen, the best model for generating targets not matched to a predefined list. Although TSE-mapped achieves higher HE score in explicit case, its BTSD score is lower. This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix K). In contrast, LLMs mostly generate targets that are highly or partially related to gold targets, resulting in fewer irrelevant outputs and higher BTSD, despite slightly lower average HE scores.
When the target is not explicitly mentioned, LLMs produce higher-quality targets than both TSE models. However, the quality drops for LLMs (TSE as well) compared to explicit cases (as per Table 2), with HE scores slightly lower than $0.5$ (refers to partially related target) on average. A manual inspection of the lower quality targets suggest that the lower score in non-explicit cases is due to insufficient context of the text. The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies. For instance, for the post "I mean, this reads to me like anyone who’s an alcoholic doesn’t go to heaven…", GPT-3.5 generated the target "Sinful behavior and heaven’s eligibility" instead of the correct gold target "gay rights".
When comparing tested proprietary and open LLMs using BTSD and HE, there is no clear advantage in TG (i.e., not all proprietary models outperform open ones, or vice versa). Overall, GPT-4o produces targets of equal or higher quality than other tested proprietary models across the approaches, including explicit and non-explicit target mentions. Among the tested open LLMs, Mistral-large and Llama-3-70B outperform each other under different settings.
Comparing between approaches, most models perform better in TG&SD across both datasets, for explicit and non-explicit cases.
@@ 5.2 Stance Detection (SD)
As shown in Table 2, according to the SC (stance classification), LLMs outperform both TSE models in detecting stances in both explicit and non-explicit cases. In general, LLMs perform better in SD when the target is explicitly mentioned in the text compared to when it is not. Our manual analysis shows that in non-explicit cases, the input text lacks sufficient context about the target, leading to lower SC scores.
The tested proprietary and open LLMs exhibit competitive performance against each other in both explicit and non-explicit cases for both approaches, with neither showing a clear advantage.
When comparing approaches, LLMs generally perform better with the holistic TG&SD strategy in both explicit and non-explicit cases, with the exception of Llama-3-8B, and Mistral-small (for non-explicit).
The quality of the generated target (BTSD) shows a stronger correlation with the SC score in explicit cases (avg. $\tau=0.4$) across datasets compared to non-explicit ones (avg. $\tau=0.25$), particularly for VAST and EZSTANCE.
Upon manual investigation, we observed that LLMs such GPT-4o may generate highly relevant targets that are closely related to the gold targets but nearly ‘antithetical’ or opposite, leading to a stance reversal in comparison to the gold stance. For example, in VAST, when the gold target is "permit to carry gun", GPT-4o generates "gun control", while GPT-3.5 generates "guns". Although GPT-4o’s target is more closely related to the gold target, it conveys the opposite. GPT-3.5’s target, though broader, aligns with the stance of gold target. Another example, in the explicit sample text, "Q. from audience: Malcom X said the most disrespected person in America is the black woman, has #MeToo changed this? Gender has been constructed in the image of the white woman, the idea of a level playing field, white women are often in better positions, says @HeidiMirza #CHevents", GPT-3.5 generated the target "Gender roles", which, while related, fails to capture the specific stance-related target "MeToo movement". This mistake reflects the model’s need for broader context to establish the correct link. A potential way to address this issue would be leveraging an LLM as a judge to assess the stance alignment. We explore the feasibility of this idea by conducting a small-scale experiment in Appendix J.
@@ 6 Conclusion
  • This work has highlighted specific challenges of the proposed OTSD task and measured both target quality and stance detection performance. However, our findings underscore the need to improve the measure of stance detection accuracy in OTSD, enabling it to account for the alignment/coherence between the generated target and the stance toward it, given the text and the reference (gold) target. To investigate the feasibility of a potential solution, we carried out a preliminary experiment using an LLM as an automated judge of stance-target coherence. The LLM’s assessments showed strong correlation with human judgments, suggesting that LLMs can serve as a reliable component in the development of coherence-based evaluation metrics. Full details of this pilot study are provided in Appendix J.
  • In this work, we focus on identifying a single target from each text, even though a given text may contain multiple targets, whether explicitly stated or implied.
@@
  • Although our findings were evaluated across a diverse set of eight LLMs spanning four different families—covering both small and large models, as well as proprietary and open-source ones—future work could benefit from exploring a broader range of models, particularly Large Reasoning Models (LRMs), which present an especially promising direction.
Ethics Statement
The datasets used for fine-tuning and zero-shot experiments are publicly available and anonymized, with no associated data protection concerns. Additionally, the human annotators involved in the evaluation were informed about the intended use of their contributions and participated voluntarily without compensation. The only data collected consisted of their annotation ratings.
Acknowledgments
This research was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) under the Grant No. RGPIN-2022-04789, and by the Digital Research Alliance of Canada (https://www.alliancecan.ca/en). We would like to thank the annotators and the reviewers for their valuable feedback.
@@ Appendix F Dataset Samples
Following TSE, our targets are Creationism (Somasundaran and Wiebe, 2010), Gay Rights (Somasundaran and Wiebe, 2010), Climate Change is a Concern (Mohammad et al., 2016a), MeToo Movement (Gautam et al., 2020), Merger of Disney and Fox (Conforti et al., 2020), and Lockdown in New York State (Miao et al., 2020). Samples from TSE and VAST in both explicit and non-explicit settings are given in Table 4 and 5, respectively. The demonstration of samples can also be found in the accompanying GitHub repository [^3] GitHub link for sample demonstration: https://github.com/AbuUbaida/opentarget..
@@ Appendix G SemSim Details
Appendix H Human Annotation Process
Three graduate students, specializing in NLP, participated voluntarily as annotators in the evaluation of generated targets. They were chosen for their expertise and familiarity with the stance detection task, enabling them to effectively assess the relevance of the generated targets compared to the golden ones. First, we provided them with an annotation guideline (detailed in §H.1) that included descriptions of the datasets, an overview of the task, and a descriptive scoring scale accompanied by examples. Each annotator then independently scored the targets generated by different LLMs for 500 randomly selected samples (300 explicit and 200 non-explicit). To assess inter-annotator agreement, we calculated Krippendorff’s $\alpha$ (Krippendorff, 2011) and Fleiss’s $\kappa$ (Fleiss, 1971), with the results presented in Table 6. The final score for each generated target was determined based on the majority vote among the annotators. In cases of a tie, the median score was used.
H.1 Annotation Guideline
@@
BTSD (GPT-3.5) 38.47 (38.43$\uparrow$ 41.10 (41.67$\downarrow$ 47.60 (49.27$\downarrow$ 37.88 (39.60$\downarrow$ 41.34 (44.25$\downarrow$ 48.98 (49.70$\downarrow$
BTSD (GPT-4o) 40.63 (41.55$\downarrow$ 41.59 (42.69$\downarrow$ 48.12 (49.70$\downarrow$ 41.31 (41.92$\downarrow$ 42.33 (44.25$\downarrow$ 48.76 (50.69$\downarrow$
SC (GPT-3.5) 37.91 (42.68$\downarrow$ 47.47 (47.21$\uparrow$ 40.12 (39.10$\uparrow$ 44.86 (47.61$\downarrow$ 47.00 (48.48$\downarrow$ 40.74 (40.63$\uparrow$
SC (GPT-4o) 47.36 (44.78$\uparrow$ 41.17 (40.77$\uparrow$ 43.79 (45.93$\downarrow$ 49.14 (46.83$\uparrow$ 47.84 (49.38$\downarrow$ 45.68 (46.22$\downarrow$
@@
BTSD (GPT-3.5) 30.73 (33.10$\downarrow$ 41.54 (38.10$\uparrow$ 40.76 (38.78$\uparrow$ 32.14 (31.32$\uparrow$ 40.38 (38.55$\uparrow$ 40.54 (41.89$\downarrow$
BTSD (GPT-4o) 32.72 (35.14$\downarrow$ 41.71 (38.55$\uparrow$ 41.34 (42.78$\downarrow$ 33.44 (36.12$\downarrow$ 41.92 (39.84$\uparrow$ 43.90 (45.70$\downarrow$
SC (GPT-3.5) 31.88 (32.66$\downarrow$ 43.41 (45.54$\downarrow$ 37.76 (36.69$\uparrow$ 35.46 (33.94$\uparrow$ 44.46 (45.80$\downarrow$ 38.10 (38.97$\downarrow$
SC (GPT-4o) 36.00 (36.39$\downarrow$ 38.27 (39.92$\downarrow$ 38.34 (38.50$\downarrow$ 38.11 (37.5$\uparrow$ 42.81 (43.84$\downarrow$ 37.54 (39.26$\downarrow$
Table 7: Preliminary results using the CoT prompting strategy with GPT-3.5 and GPT-4o models on three datasets (TSE, VAST, EZSTANCE), evaluated for both Target Generation (TG) using BTSD and Stance Detection (SD) using SC. Results in parentheses refer to scores reported in Table 2 using the "Task Definition" prompting strategy. An upward arrow ($\uparrow$) indicates that the CoT strategy outperformed the Task Definition approach, while a downward arrow ($\downarrow$) indicates a performance drop. The results are presented for both Explicit (Ex) and Non-explicit (N-Ex) cases across two approaches: TG+SD and TG&SD.
Appendix I Chain-of-Thought (CoT) Results
@@
Appendix J Preliminary Study on Coherence Measure in OTSD
To address the need for a more accurate evaluation metric in the Open Target Stance Detection (OTSD) task—one that aligns the generated stance with the corresponding generated target—we conduct a small-scale experiment that might be leading to a promising direction for developing a robust coherence-based measure. Specifically, we employ an LLM to assess how coherent a generated stance is with respect to the generated target, given the context of a text.
For our experiment, we selected 200 samples from the EZSTANCE dataset (Zhao and Caragea, 2024), covering all configurations (i.e., Explicit and Non-explicit, TG+SD and TG&SD). We use LLaMA-3-8B as the evaluation model, and choose the same "Task Definition" prompting strategy as applied in Section 3.2 but with few-shot samples. Each prompt includes the input text, the generated target, and the corresponding generated stance. As this is a preliminary experiment, future work could explore a variety of alternative prompting strategies to further improve performance. The prompt used in our work is as follows:
"You will be given a text, a generated stance (generated_stance), and a generated target (generated_target). Your task is to determine whether the stance is coherent with respect to the target in the context of the text. A stance is coherent if it correctly reflects the text’s position towards the target as implied or stated in the text. Analyze the text well, as it may convey the stance towards the target implicitly. Output 1 if the stance is coherent, and 0 if it is not. Output only a single digit: 0 or 1, with no explanation.
Input Format:
text: <tweet text>
generated_target: <target>
generated_stance: <stance>
Examples:
text: "Renewable energy is the only viable path forward if we care about this planet."
generated_target: renewable energy
generated_stance: AGAINST
Output: 0
text: "Elon Musk is brilliant, but his take on public transport is just awful."
generated_target: Elon Musk
generated_stance: FAVOR
Output: 1
text: "Vaccines have done more good than harm. I trust the science."
generated_target: vaccines
generated_stance: NONE
Output: 0
text: "As you have been told already what Phil says is irrelevant to ID. Or you have to allow what Dawkins says about the theory of evolution.What part of that don’t you understand?And it is a fact that if you could support your position with real data ID would go away. You can whine all you want but that will not change that fact."
generated_target: Intelligent Design
generated_stance: AGAINST
Output: 1
text: "I’ve seen hundreds of tweets at this point, many from people with big platforms, condemning leftists who are apparently pro-Putin. I have yet to see any leftists actually express support for Putin. All I’ve seen from the left is unambiguous solidarity for the people of Ukraine."
generated_target: leftists
generated_stance: NONE
Output: 1"
To further validate the LLM-based evaluation, we also collected human judgments on the same 200 samples. A single human annotator was asked to assess the coherence of the generated stances with respect to the generated targets, using the same task definition provided to the LLM. The annotator received the same input triplet: text, generated target, and generated stance.
We then measured the correlation between the LLM’s coherence judgments and those of the human annotator using the Phi coefficient. The resulting score was 0.6541, indicating a strong positive correlation (Njuka and Phiri, 2021).
It is important to note that we do not propose a definitive solution or formal methodology for coherence evaluation in stance detection. Rather, we present this as a potential path that can be further explored by the research community. Our experiment suggests that LLMs might serve as reliable judges for assessing the coherence between a generated stance and target within a given text, with judgments that align well with human evaluations. To strengthen the validity of this experiment, future studies could incorporate a wider range of LLMs, multiple runs per sample, and a larger, more diverse set of samples across the three datasets.
@@
Figure 1: Distribution of human evaluation scores for target relevance assessment (0 – Not Related, 0.5 – Partially or Indirectly Related, 1 – Completely Related). The final score for each sample is determined by the majority vote among the three annotators. In cases of a tie, the average score of 0.5 is assigned. The details on human annotator guideline are provided in Appendix H.1, and the annotator agreement is in Appendix Table 6.